What is the orbital designation used by chemists (e.g. 1s, 3d, etc.) to represent the atomic orbital with two units of orbital angular momentum and a Radial function that has zeroes at 4 different values of r?
Answer #1

orbital angular momentum has 2 values means orbital is p (0, = or -1)

An orbital has n-1 radial nodes. That is n= 5 (four radial nodes are present)

orbital is 5p

part B

one unit for angular momentum is orbital is s (m=0)

3 rasial nodes means n= 3+1 =4

orbital is 4s